# EraseDiff: Erasing Data Influence in Diffusion Models

### v3

@@ Abstract
We introduce EraseDiff, an unlearning algorithm designed for diffusion models to address concerns related to data memorization. Our approach formulates the unlearning task as a constrained optimization problem, aiming to preserve the utility of the diffusion model on retained data while removing the information associated with the data to be forgotten. This is achieved by altering the generative process to deviate away from the ground-truth denoising procedure. To manage the computational complexity inherent in the diffusion process, we develop a first-order method for solving the optimization problem, which has shown empirical benefits. Extensive experiments and thorough comparisons with state-of-the-art algorithms demonstrate that EraseDiff effectively preserves the model’s utility, efficacy, and efficiency.
@@
Figure 1: Top to Bottom: generated samples by SD v1.4 and model scrubbed by our method, EraseDiff, when erasing the concept of ‘nudity’. EraseDiff can avoid NSFW content while preserving model utility. Source code is available at https://github.com/JingWu321/EraseDiff.
@@ 1 Introduction
Diffusion Models (Ho et al., 2020; Song et al., 2020; Rombach et al., 2022) are now the method of choice in deep generative models, owing to their high-quality output, stability, and ease of training procedure. This has facilitated their successful integration into commercial applications such as midjourney. Unfortunately, the ease of use associated with diffusion models brings forth significant privacy risks. Studies have shown that these models can memorize and regenerate individual images from their training datasets (Somepalli et al., 2023a, b; Carlini et al., 2023). Beyond privacy, diffusion models are susceptible to misuse and can generate inappropriate digital content (Rando et al., 2022; Salman et al., 2023; Schramowski et al., 2023). They are also vulnerable to poison attacks (Chen et al., 2023b), allowing the generation of target images with specific triggers. These factors collectively pose substantial security threats. Moreover, the ability of diffusion models to emulate distinct artistic styles (Shan et al., 2023; Gandikota et al., 2023a) raises questions about data ownership and compliance with intellectual property and copyright laws.
@@
In this context, individuals whose images are used for training might request the removal of their private data. In particular, data protection regulations like the European Union General Data Protection Regulation (GDPR) (Voigt & Von dem Bussche, 2017) and the California Consumer Privacy Act (CCPA) (Goldman, 2020) grant users the right to be forgotten, obligating companies to expunge data pertaining to a user upon receiving a request for deletion. These legal provisions grant data owners the right to remove their data from trained models and eliminate its influence on said models (Bourtoule et al., 2021; Guo et al., 2020; Golatkar et al., 2020; Mehta et al., 2022; Sekhari et al., 2021; Ye et al., 2022; Tarun et al., 2023b, a; Chen et al., 2023a).
@@
A straightforward solution for unlearning is to retrain the model from scratch after excluding the data that needs to be forgotten. However, the removal of pertinent data followed by retraining diffusion models from scratch demands substantial resources and is often deemed impractical. A version of the stable diffusion model trained on subsets of the LAION-5B dataset (Schuhmann et al., 2022) costs approximately 150,000 GPU hours with 256 A100 GPUs [^1] https://stablediffusion.gitbook.io/overview/stable-diffusion-overview/technology/training-procedures. Existing research on efficient unlearning have primarily focused on classification problems (Karasuyama & Takeuchi, 2010; Cao & Yang, 2015; Ginart et al., 2019; Bourtoule et al., 2021; Wu et al., 2020; Guo et al., 2020; Golatkar et al., 2020; Mehta et al., 2022; Sekhari et al., 2021; Chen et al., 2023a). Despite substantial progress, methods developed for unlearning in classification are observed to be ineffective for generation tasks as studied by Fan et al. (2023). Consequently, there is a pressing need for the development of methods capable of scrubbing data from diffusion models without necessitating complete retraining.
Recently, a handful of studies (Gandikota et al., 2023a, b; Zhang et al., 2023; Heng & Soh, 2023a, b; Fan et al., 2023) target unlearning in diffusion models, with a primary focus on the text-to-image models (Gandikota et al., 2023a, b; Zhang et al., 2023; Bui et al., 2024). Heng & Soh (2023b) utilize ideas from continual learning to preserve model utility when performing forgetting for a wide range of generative models. Their method requires the computation of the Fisher Information Matrix (FIM) for different datasets and models, which could lead to significant computational demands. Fan et al. (2023) propose to shift the attention to salient weights w.r.t. the forgetting data, resulting in a very potent unlearning algorithm across image classification and generation tasks.
In this work, we propose EraseDiff, and formulate diffusion unlearning as a constrained Optimization problem, where the objective is to finetune the models with the remaining data ${\mathcal{D}}_{r}$ for preserving the model utility and to erase the influence of the forgetting data ${\mathcal{D}}_{f}$ on the models by deviating the learnable reverse process from the ground-truth denoising procedure, namely minimizing the loss over the remaining data while maximizing that over the forgetting data. A common issue in unlearning is the gradient conflict, as optimizing one objective could hinder another one. To address this issue, we adopt an approximate optimization problem that identifies an optimal direction to update different objectives. We benchmark EraseDiff on various scenarios, encompassing unlearning of classes on CIFAR10 (Krizhevsky et al., 2009) with Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020), classes on Imagenette (Howard & Gugger, 2020) and concepts on the I2P dataset (Schramowski et al., 2023) with stable diffusion. Our empirical findings show that EraseDiff is $\times1$ faster than Heng and Soh’s method (Heng & Soh, 2023b) and 2×2$\times$2 × faster than Fan’s method (Fan et al., 2023) when forgetting on DDPM while achieving better unlearning results across several metrics. The results demonstrate that EraseDiff is capable of effectively erasing data influence in diffusion models, ranging from specific classes to the concept of nudity.
@@ 2 Background
In this section, we outline the components of the models we evaluate, including DDPM, classifier-free guidance diffusion models (Ho & Salimans, 2022), Latent Diffusion Models (LDM) (Rombach et al., 2022). Throughout the paper, we denote scalars, and vectors/matrices by lowercase and bold symbols, respectively (e.g., $a$, ${\bm{a}}$, ${\bm{A}}$).
@@
DDPM. (1) Diffusion: DDPM gradually diffuses the data distribution ${\mathbb{R}}^{d}\ni{\mathbf{x}}_{0}\sim q({\mathbf{x}})$ into the standard Gaussian distribution ${\mathbb{R}}^{d}\ni{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},\mathbf{I}_{d})$ with $T$ time steps, ie., $q({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1})={\mathcal{N}}({\mathbf{x}}_{t};\sqrt{\alpha_{t}}{\mathbf{x}}_{t-1},(1-\alpha_{t})\mathbf{I}_{d})$, where $\alpha_{t}=1-\beta_{t}$ and $\{\beta_{t}\}^{T}_{t=1}$ are the pre-defined variance schedule. The diffusion takes the form ${\mathbf{x}}_{t}$ as ${\mathbf{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$, where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. (2) Training: A model $\epsilon_{{\bm{\theta}}}(\cdot)$ with parameters ${\bm{\theta}}\in{\mathbb{R}}^{n}$ is trained to learn the reverse process $p_{{\bm{\theta}}}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})\approx q({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$. Given ${\mathbf{x}}_{0}\sim q({\mathbf{x}})$ and time step $t\in[1,T]$, the simplified training objective is to minimize the distance between ${\bm{\epsilon}}$ and the predicted ${\bm{\epsilon}}_{t}$ given ${\mathbf{x}}_{0}$ at time $t$, ie., $\|{\bm{\epsilon}}-\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t},t)\|$. (3) Sampling: after training the model, we could obtain the learnable backward distribution $p_{{\bm{\theta}}^{*}}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})={\mathcal{N}}({\mathbf{x}}_{t-1};{\bm{\mu}}_{{\bm{\theta}}^{*}}({\mathbf{x}}_{t},t),{\bm{\Sigma}}_{{\bm{\theta}}^{*}}({\mathbf{x}}_{t},t))$, where ${\bm{\mu}}_{{\bm{\theta}}^{*}}({\mathbf{x}}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}({\mathbf{x}}_{t}-\frac{\beta_{t}}{\sqrt{1-\alpha_{t}}}\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t},t))$ and ${\bm{\Sigma}}_{{\bm{\theta}}^{*}}({\mathbf{x}}_{t},t)=\frac{(1-\bar{\alpha}_{t-1})\beta_{t}}{1-\bar{\alpha}_{t}}$. Then, given ${\mathbf{x}}_{T}\sim{\mathcal{N}}({\bm{0}},\mathbf{I}_{d})$, ${\mathbf{x}}_{0}$ could be obtained via sampling from $p_{{\bm{\theta}}^{*}}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ from $t=T$ to $t=1$ step by step.
Classifier-free guidance. Classifier-free guidance is a conditioning method to guide diffusion-based generative models without an external pre-trained classifier. Model prediction would be $\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t}|c,t)$, where $c$ is the input’s corresponding label. The unconditional and conditional models are jointly trained by randomly setting $c$ to the unconditional class $\emptyset$ identifier with the probability $p_{uncond}$. Then, the sampling procedure uses the linear combination of the conditional and unconditional score estimates as ${\bm{\epsilon}}_{t}=(1+w)\cdot\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t}|c)-w\cdot\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t}|\emptyset)$, $w$ is the guidance scale that controls the strength of the classifier guidance.
@@
Latent diffusion model. Latent diffusion models (LDM) apply the diffusion models in the latent space ${\mathbf{z}}$ of a pre-trained variational autoencoder. The noise would be added to ${\mathbf{z}}=\varepsilon({\mathbf{x}})$, instead of the data ${\mathbf{x}}$, and the denoised output would be transformed to image space with the decoder. Besides, cross-attention layers are introduced into the model for general conditioning inputs.
@@ 3.1 Training objective
${\mathcal{L}}_{f}({\bm{\theta}};{\mathcal{D}}_{f})=\mathbb{E}_{t,{\bm{\epsilon}}\in{\mathcal{N}}({\bm{0}},\mathbf{I}_{d}),({\mathbf{x}}_{0},c)\sim{\mathcal{D}}_{f}\times{\mathcal{C}}_{f}}[\|{\bm{\epsilon}}_{f}-\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t}|c)\|_{2}^{2}].$ (2)
With this, we hinder the approximator $\epsilon_{{\bm{\theta}}}$ to guide the denoising process to obtain meaningful examples for the forgetting data example ${\mathbf{x}}_{0}\sim{\mathcal{D}}_{f}$. In our experiments, we choose ${\bm{\epsilon}}_{f}$ to be a distribution different from ${\bm{\epsilon}}\in{\mathcal{N}}({\bm{0}},\mathbf{I}_{d})$. This could be ${\bm{\epsilon}}_{f}\in{\mathcal{U}}({\bm{0}},\mathbf{I}_{d})$, or $\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t}|c_{m})$ like Fan et al. (2023); Heng & Soh (2023b) where $c_{m}\neq c$ so that the denoised image ${\mathbf{x}}_{0}$ is not related to the forgetting class/concept $c$.
To perform unlearning and minimize ${\mathcal{L}}_{r}({\bm{\theta}};{\mathcal{D}}_{r})$ and ${\mathcal{L}}_{f}({\bm{\theta}};{\mathcal{D}}_{f})$ simultaneously, it is common to form
@@
with $\lambda\geq 0$ as the optimization objective (see for example Fan et al. (2023)). However, training could be hindered due to the conflicting gradients between the retaining and forgetting objectives.
Equation 3 could also be viewed as a scalarization of a Multi-Objective Optimization (MOO) problem, ie., minimizing $\big{(}{\mathcal{L}}_{r}({\bm{\theta}};{\mathcal{D}}_{r}),{\mathcal{L}}_{f}({\bm{\theta}};{\mathcal{D}}_{f})\big{)}^{\top}$. It is well known that MOO should address the gradient conflict issue.
Instead of scalarization of MOO, we propose to minimize the following objective:
@@
s.t. $\operatorname{min}_{\bm{\phi}}{\mathcal{L}}_{f}(\bm{\phi};{\mathcal{D}}_{f},\bm{\phi}_{\texttt{init}}={\bm{\theta}})\;.$ (4)
Here, the problem $\operatorname{min}_{\bm{\phi}}\;{\mathcal{L}}_{f}(\bm{\phi};{\mathcal{D}}_{f},\bm{\phi}_{\texttt{init}}={\bm{\theta}})$ indicates that given ${\bm{\theta}}$, the optimization of $\bm{\phi}$ starts from ${\bm{\theta}}$ and aims to minimize the forgetting loss. In other words, if optimality ${\bm{\theta}}^{\ast}$ is achieved, we have found ${\bm{\theta}}^{\ast}$ that maintains the model’s utility as a result of $\operatorname{min}_{{\bm{\theta}}}\;{\mathcal{L}}_{r}({\bm{\theta}};{\mathcal{D}}_{r})$, and starting from ${\bm{\theta}}^{\ast}$, we cannot further reduce the forgetting loss due to $\operatorname{min}_{\bm{\phi}}\;{\mathcal{L}}_{f}(\bm{\phi};{\mathcal{D}}_{f},\bm{\phi}_{\texttt{init}}={\bm{\theta}}^{\ast})$. Now, we note that $\operatorname*{arg\,min}_{{\bm{\theta}}}\left\{{\mathcal{L}}({\bm{\theta}})-\operatorname{min}_{{\bm{\theta}}}{\mathcal{L}}({\bm{\theta}})\leq 0\right\}\in\{\operatorname*{arg\,min}_{{\bm{\theta}}}{\mathcal{L}}({\bm{\theta}})\}.$ This insight will aid us in solving Section 3.1 efficiently as we will show next. Putting everything together, we propose:
@@
s.t. ${\mathcal{L}}_{f}({\bm{\theta}};{\mathcal{D}}_{f})-\operatorname{min}_{\bm{\phi}}{\mathcal{L}}_{f}(\bm{\phi};{\mathcal{D}}_{f},\bm{\phi}_{\texttt{init}}={\bm{\theta}})\leq 0,$ (5)
where $\bm{\phi}$ is initialized at ${\bm{\theta}}$.

### v4

# Erasing Undesirable Influence in Diffusion Models

@@ Abstract
Diffusion models are highly effective at generating high-quality images but pose risks, such as the unintentional generation of NSFW (not safe for work) content. Although various techniques have been proposed to mitigate unwanted influences in diffusion models while preserving overall performance, achieving a balance between these goals remains challenging. In this work, we introduce EraseDiff, an algorithm designed to preserve the utility of the diffusion model on retained data while removing the unwanted information associated with the data to be forgotten. Our approach formulates this task as a constrained optimization problem using the value function, resulting in a natural first-order algorithm for solving the optimization problem. By altering the generative process to deviate away from the ground-truth denoising trajectory, we update parameters for preservation while controlling constraint reduction to ensure effective erasure, striking an optimal trade-off. Extensive experiments and thorough comparisons with state-of-the-art algorithms demonstrate that EraseDiff effectively preserves the model’s utility, efficacy, and efficiency.
@@
Figure 1: Top to Bottom: generated samples by SD v1.4 and model scrubbed by our method, when erasing the concept of ‘nudity’. Our method can avoid NSFW (not safe for work) content while preserving model utility.
@@ 1 Introduction
Diffusion Models [30, 59, 48] are now the method of choice in deep generative models, owing to their high-quality output, stability, and ease of training procedure. This has facilitated their successful integration into commercial applications such as midjourney. Unfortunately, the ease of use associated with diffusion models brings forth significant privacy risks. Studies have shown that these models can memorize and regenerate individual images from their training datasets [57, 58, 10]. Beyond privacy, diffusion models are susceptible to misuse and can generate NSFW digital content [47, 51, 53]. In this context, individuals whose images are used for training might request the removal of their private data. In particular, data protection regulations like the European Union General Data Protection Regulation (GDPR) [63] and the California Consumer Privacy Act (CCPA) [24] grant users the right to be forgotten, obligating companies to expunge data pertaining to a user upon receiving a request for deletion. These legal provisions grant data owners the right to remove their data from trained models and eliminate its influence on said models [4, 25, 23, 43, 55, 67, 60, 61, 11].
@@
A straightforward solution is to retrain the model from scratch after excluding the data that needs to be forgotten. However, the removal of pertinent data followed by retraining diffusion models from scratch demands substantial resources and is often deemed impractical. A version of the stable diffusion model trained on subsets of the LAION-5B dataset [54] costs approximately 150,000 GPU hours with 256 A100 GPUs [^1] https://stablediffusion.gitbook.io/overview/stable-diffusion-overview/technology/training-procedures. Existing research on erasing unwanted influence has primarily focused on classification tasks [33, 7, 22, 4, 66, 25, 23, 43, 55, 11]. Despite substantial progress, prior methods developed in classification are observed to be ineffective for generation tasks [16]. Consequently, there is a pressing need for the development of methods capable of scrubbing data from diffusion models without necessitating complete retraining.
Recently, a handful of studies [20, 21, 69, 27, 28, 35, 16, 40, 6] target unlearning in diffusion models, with a primary focus on the text-to-image models [20, 21, 69]. Broadly, these methods aim to achieve two main objectives: erasing data influence and preserving overall model performance. However, as demonstrated by Bui et al. [5], balancing this trade-off remains challenging.
In this work, we propose EraseDiff, an algorithm tailored to balance the overall performance of diffusion models with the erasure of undesirable information. Drawing inspiration from optimization-based meta-learning algorithms [41, 18] that enable fast adaptation to new learning tasks, we formulate this challenge as a bi-level optimization problem, where the “inner” optimization focus on erasing undesirable influence and the “outer” objective seeks to preserve model performance. The outer objective and inner optimization are interdependent, iterating between preservation and erasure to balance the trade-off effectively. However, this nested optimization can be challenging to optimize efficiently. The inner optimization may converge to a saddle point or struggle with non-convex functions, making it difficult to achieve a stable solution [38]. Therefore, we further reformulate the problem as a constrained optimization problem using the value function [39, 45, 68], which facilitates a natural first-order solution [38], allows us to optimize preservation and erasure in a unified manner. This approach achieves a fine-tuned balance between preservation and targeted erasure, yielding an optimal trade-off. We benchmark EraseDiff on various scenarios, encompassing unlearning of classes on CIFAR-10 [34] with Denoising Diffusion Probabilistic Models (DDPM) [30], classes on Imagenette [31] and concepts on the I2P dataset [53] with stable diffusion. Our empirical findings show that EraseDiff is $\times1$ faster than Heng and Soh’s method [28] and 2×2$\times$2 × faster than Fan’s method [16] when forgetting on DDPM while achieving better unlearning results across several metrics. The results demonstrate that EraseDiff is capable of effectively erasing data influence in diffusion models, ranging from specific classes to the concept of nudity.
@@ 2 Background
In this section, we outline the components of the models we evaluate, including DDPM and latent diffusion models [48]. Throughout the paper, we denote scalars, and vectors/matrices by lowercase and bold symbols, respectively (e.g., $a$, ${\bm{a}}$, ${\bm{A}}$).
DDPM.
@@
(1) Diffusion: DDPM gradually diffuses the data distribution ${\mathbb{R}}^{d}\ni{\mathbf{x}}_{0}\sim q({\mathbf{x}})$ into the standard Gaussian distribution ${\mathbb{R}}^{d}\ni{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},\mathbf{I}_{d})$ with $T$ time steps, i.e., $q({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1})={\mathcal{N}}({\mathbf{x}}_{t};\sqrt{\alpha_{t}}{\mathbf{x}}_{t-1},(1-\alpha_{t})\mathbf{I}_{d})$, where $\alpha_{t}=1-\beta_{t}$ and $\{\beta_{t}\}^{T}_{t=1}$ are the pre-defined variance schedule. The diffusion takes the form ${\mathbf{x}}_{t}$ as ${\mathbf{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$, where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. (2) Training: A model $\epsilon_{{\bm{\theta}}}(\cdot)$ with parameters ${\bm{\theta}}\in{\mathbb{R}}^{n}$ is trained to learn the reverse process $p_{{\bm{\theta}}}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})\approx q({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$. Given ${\mathbf{x}}_{0}\sim q({\mathbf{x}})$ and time step $t\in[1,T]$, the simplified training objective is to minimize the distance between ${\bm{\epsilon}}$ and the predicted ${\bm{\epsilon}}_{t}$ given ${\mathbf{x}}_{0}$ at time $t$, i.e., $\|{\bm{\epsilon}}-\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t},t)\|$. (3) Sampling: after training the model, we could obtain the learnable backward distribution $p_{{\bm{\theta}}^{*}}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})={\mathcal{N}}({\mathbf{x}}_{t-1};{\bm{\mu}}_{{\bm{\theta}}^{*}}({\mathbf{x}}_{t},t),{\bm{\Sigma}}_{{\bm{\theta}}^{*}}({\mathbf{x}}_{t},t))$, where ${\bm{\mu}}_{{\bm{\theta}}^{*}}({\mathbf{x}}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}({\mathbf{x}}_{t}-\frac{\beta_{t}}{\sqrt{1-\alpha_{t}}}\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t},t))$ and ${\bm{\Sigma}}_{{\bm{\theta}}^{*}}({\mathbf{x}}_{t},t)=\frac{(1-\bar{\alpha}_{t-1})\beta_{t}}{1-\bar{\alpha}_{t}}$. Then, given ${\mathbf{x}}_{T}\sim{\mathcal{N}}({\bm{0}},\mathbf{I}_{d})$, ${\mathbf{x}}_{0}$ could be obtained via sampling from $p_{{\bm{\theta}}^{*}}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ from $t=T$ to $t=1$ step by step.
Latent diffusion models.
@@
Latent diffusion models apply the diffusion models in the latent space ${\mathbf{z}}$ of a pre-trained variational autoencoder. The noise would be added to ${\mathbf{z}}=\varepsilon({\mathbf{x}})$, instead of the data ${\mathbf{x}}$, and the denoised output would be transformed to image space with the decoder. Besides, text embeddings generated by models like CLIP are used as conditioning inputs.
@@ 3.1 Training objective
${\mathcal{L}}_{f}({\bm{\theta}};{\mathcal{D}}_{f})=\mathbb{E}_{t,{\bm{\epsilon}}\in{\mathcal{N}}({\bm{0}},\mathbf{I}_{d}),({\mathbf{x}}_{0},c)\sim{\mathcal{D}}_{f}\times{\mathcal{C}}_{f}}[\|{\bm{\epsilon}}_{f}-\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t}|c)\|_{2}^{2}],$ (2)
where ${\bm{\epsilon}}_{f}=\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t}|c_{m})$ and $c_{m}\neq c$ so that the denoised image ${\mathbf{x}}_{0}$ is not related to the forgetting class/concept $c$ [16, 28]. With this, we hinder the approximator $\epsilon_{{\bm{\theta}}}$ to guide the denoising process to obtain meaningful examples for the forgetting data example ${\mathbf{x}}_{0}\sim{\mathcal{D}}_{f}$.
To erase the undesirable influence of ${\mathcal{D}}_{f}$ and preserve the overall performance, it is common to form
@@
with $\lambda>0$ as the optimization objective (see for example [16]). However, training could be hindered due to the conflicting gradients between the erasing and preservation objectives, preventing a balanced trade-off between erasure and preservation [37]. To address this, rather than scalarizing the two objectives, we consider a framework based on optimization-based meta-learning algorithm [46] that allows iteratively updates to optimize each objective:
@@
s.t. ${\bm{\theta}}\in\operatorname*{arg\,min}_{\bm{\phi}}{\mathcal{L}}_{f}(\bm{\phi};{\mathcal{D}}_{f})\;,$ (4)
where the outer objective minimizes the remaining loss ${\mathcal{L}}_{r}$ (i.e., preserving model utility), the inner optimization minimizes the forgetting loss ${\mathcal{L}}_{f}$ (i.e., erasing) with initialization $\mathbb{R}^{n}\ni\bm{\phi}_{\texttt{init}}={\bm{\theta}}$. Given ${\bm{\theta}}$, the inner optimization on $\bm{\phi}$ aims to minimize the forgetting data influence, with the goal of achieving effective erasure while preserving model utility. The outer objective and inner optimization are interdependent, iterating between preservation and erasure to balance the trade-off effectively.
While the above framework allows for iterative updates to address the conflicting objectives of erasure and preservation, it still relies on nested optimization, which can be challenging to optimize efficiently. The inner optimization may converge to a saddle point or struggle with non-convex functions, making it difficult to achieve a stable solution [38]. To further streamline the optimization process, we adopt a value function approach [39, 45, 68] that reformulates the problem as a single-constrained optimization:
@@
s.t. ${\mathcal{L}}_{f}({\bm{\theta}};{\mathcal{D}}_{f})-\operatorname{min}_{\bm{\phi}}{\mathcal{L}}_{f}(\bm{\phi};{\mathcal{D}}_{f})\leq 0,$ (5)
where $\bm{\phi}$ is initialized at ${\bm{\theta}}$, leverages the value function to encapsulate the influence of data erasure directly as a constraint on ${\mathcal{L}}_{f}$. This avoids the need for a nested loop by capturing the forgetting objective as a constraint and provides a natural first-order solution, as the constrained formulation allows us to optimize preservation and erasure in a unified manner.